I can not do this problem, and anyone help me please?

Solve the following system by graphing:
3x - 6y = 9
x - 2y = 3
In EQUATION2 show:y-intercept as an ordered pair =>
x-intercept as an ordered pair =>
What is the solution for this system?

What type of system is this?

Answer by checkley71(8403)

USE THE Y=mX+b METHOD TO FIND THE SLOPE AND THE X & Y INTERCEPTS PLOT THE TWO LINES AND DEFINE THE INTERSECTION.
THUS 3X-6Y=9 OR -6Y=-3X+9 OR Y=X/2-3/2 THAT GIVES YOU A SLOPE OF 1/2 & A Y INTERCEPT OF -3/2. THE X VALUE IS OBTAINED BY SETTING Y=0 0=X/2-3/2 OR X/2=3/2 OR X=3 SO THESE TWO OINTS ARE (0,-3/2) & (3,0)
THE SECOND EQUATION IS SOLVED THE SAME WAY X-2Y=3 OR -2Y=-X+3 OR Y=X/2-3/2
WHAT THIS TELLS YOU IS THAT THESE TWO EQUATIONS ARE FOR THE SAME LINE HAVING THE SAME SLOPE, THE SAME Y INTERCEPT AND THE SAME TWO POINTS.